Can I know the formula for work done in an adiabatic process in terms of only pressure ?

Grade:12

3 Answers

Arun
6 years ago
Dear Shubham
 
  • For an adiabatic process of ideal gas equation we have
         PVγ = K (Constant)               
    Where γ is the ratio of specific heat (ordinary or molar) at constant pressure and at constant voluume
         γ = Cp/Cv
  • Suppose in an adiabatic process pressure and volume of a sample of gas changs from (P1, V1) to (P2, V2) then we have
         P1(V1)γ=P2(V2)γ=K 
    Thus, P = K/Vγ
  • Work done by gas in this process is
         W = ∫PdV
    where limits of integration goes from V1 to V2
    Putting for P=K/Vγ, and integrating we get,
         W = (P1V1-P2V2)/(γ-1)         
  • In and adiabatic process if W>0 i.e., work is done by the gas then T21
  • If work is done on the gas (W2 > T1 i.e., temperature of gas rises.
Regards
Arun (askIITians forum expert)
praveen kumar
6 years ago
you use this as well  W = (P1V1-P2V2)/(γ-1)  or   the work done in the adibatic process is  
and we have another formula for work done in adibatic process in terms of temperature.     W = (P2V2-P1V1)/(1-γ) 
Anjali kumari
5 years ago
Formula of work done by adiabatic process is W=nRT.where , n is number of moles.Which is equals to W= (p2v2-p1v1)/(1-y).